By using the formula logba = (log a)/(log b) where the b is lowered solve log2 10000 where the two is lowered
Thank you in advance for your help.

Answer by mukhopadhyay(490)

log (base2) 10000 = (log 10000)/(log 2) = (log 10^4)/(log 2) = 4/(log 2)